OpenStudy (ayyookyndall):

**PLEASE HELP** Verify that parallelogram ABCD with vertices A(-5, -1), B(-9, 6), C(-1,5), and D(3,-2) is a rhombus by showing that it is a parallelogram with perpendicular diagonals.

OpenStudy (ayyookyndall):

@jim_thompson5910

OpenStudy (alexandervonhumboldt2):

use distamce formula sqrt((x1-x2)^2+(y1-y2)^2) to count the dstance. The wrte eqaution of diagonals and verify thats their slopes are negative reciprocals

OpenStudy (asnaseer):

You need to show that the slope of the line AC multiplied by the slope of the line BD is equal to -1

OpenStudy (asnaseer):

two lines are perpendicular if the product of their slopes equals -1

OpenStudy (ayyookyndall):

Okay... So now im more confused...

jimthompson5910 (jim_thompson5910):

A(-5, -1), C(-1,5) what is the slope of the line through A and C

jimthompson5910 (jim_thompson5910):

use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (ayyookyndall):

4, -6?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Slope of the line through point A and point C \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[\Large m = \frac{5-(-1)}{-1-(-5)}\] \[\Large m = \frac{5+1}{-1+5}\] \[\Large m = \frac{6}{4}\] \[\Large m = \frac{3}{2}\]

jimthompson5910 (jim_thompson5910):

I made (x1,y1) be point A (-5,-1) I made (x2,y2) be point B (-1,5)

jimthompson5910 (jim_thompson5910):

hopefully it makes sense?

OpenStudy (ayyookyndall):

So what do i say for the answer? (x1,y1) be point A (-5,-1) (x2,y2) be point B (-1,5) ???

OpenStudy (ayyookyndall):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

we want to find the slope of line AC that's why I made A the point (x1,y1) and C the point (x2,y2)

jimthompson5910 (jim_thompson5910):

I'm not dealing with point B right now

jimthompson5910 (jim_thompson5910):

AC is one diagonal BD is the other diagonal (we'll focus on that later)

OpenStudy (ayyookyndall):

Okay.. So what are we focusing on right now?

jimthompson5910 (jim_thompson5910):

finding the slope of AC

jimthompson5910 (jim_thompson5910):

I want to make sure you understand

jimthompson5910 (jim_thompson5910):

asnaseer has a good pic drawn above

jimthompson5910 (jim_thompson5910):